# Multi-Granularity Distribution Modeling for Video Watch Time Prediction via Exponential-Gaussian Mixture Network

### v1

@@ 1. INTRODUCTION
Short-video platforms such as TikTok and KuaiShou have experienced a significant surge in popularity over recent years (Savic, 2021; Wang et al., 2019a; Cai et al., 2023; Pan et al., 2023). These services typically employ a continuous swipe-based streaming play pattern without the need to click through multiple candidates explicitly (Gao et al., 2022b; Gong et al., 2022). Within this paradigm, watch time has emerged as a fundamental indicator of user satisfaction. Therefore, accurate watch time prediction is crucial (Wu et al., 2018; Yi et al., 2014; Wang et al., 2020), as it empowers the platform to recommend videos that better align with user interests and ultimately enhance user engagement.
The watch time prediction is fundamentally a regression problem since its values are continuous and span a wide range. It is well-known that the distribution of labels significantly influences the difficulty of the regression task, where appropriate distribution assumptions can enhance regression accuracy (as demonstrated by extensive existing studies (Ferrari and Cribari-Neto, 2004; Ouma et al., 2016; Wang et al., 2019b; Mardalena et al., 2020; Shono, 2008)). However, the distribution of short-video watch time remains insufficiently explored. To bridge this knowledge gap, we thoroughly investigate the distribution of short-video watch time using our real-world industrial recommendation data. As shown in Figure 1, we analyze and compare watch time distributions across coarse-to-fine-grained levels:
@@
To mitigate the two challenges mentioned above, and inspired by previous works in other fields (Ferrari and Cribari-Neto, 2004; Ouma et al., 2016; Wang et al., 2019b), we propose to directly regress the absolute values of watch time based on a sound distribution assumption, thereby improving estimation accuracy. To reconcile distributional discrepancies across different granularities, we assume that short-video watch time is governed by a latent mixture model, formally defined as the Exponential-Gaussian Mixture distribution (EGM distribution), where the exponential component addresses coarse-grained distribution skewness and the Gaussian components adaptively capture fine-grained distribution diversity.
@@ 4.2. Main Results
Watch
Time
Video
Views
Engage
Actions
MAE XAUC KL
@@ 4.2.1. Online A/B tests(RQ1)
To further demonstrate the effectiveness of EGMN, we conducted online A/B tests for 7 days on the industrial short video recommender system of Xiaohongshu App. We simultaneously attribute 10% real traffic to EGMN and our existing baseline CREAD. Our platform serves more than 150 million DAUs. Thus each traffic group covers 15 million real users and ensures the reliability of the experimental results. EGMN and CREAD are applied to the ranking stage whose predicting backbone is MMOE (Ma et al., 2018; Tang et al., 2020). We show the results of online A/B testing in the left part of Table 2. It is apparent that the proposed EGMN model achieves a significant improvement (+0.681%) in user watch time compared to the CREAD, while exhibiting no negative impact on other metrics. In addition, it is observed that EGMN also demonstrates a significant improvement (+0.189%) in Video Views, which can be attributed to its better alignment with user interests and ultimate enhancement of user engagement.

### v2

@@ 1. INTRODUCTION
Short-video platforms such as TikTok and KuaiShou have experienced a significant surge in popularity over recent years (Savic, 2021; Wang et al., 2019b; Cai et al., 2023; Pan et al., 2023). These services typically employ a continuous swipe-based streaming play pattern without the need to click through multiple candidates explicitly (Gao et al., 2022b; Gong et al., 2022). Within this paradigm, watch time has emerged as a fundamental indicator of user satisfaction. Therefore, accurate watch time prediction is crucial (Wu et al., 2018; Yi et al., 2014; Wang et al., 2020), as it empowers the platform to recommend videos that better align with user interests and ultimately enhance user engagement.
The watch time prediction is fundamentally a regression problem since its values are continuous and span a wide range. It is well-known that the distribution of labels significantly influences the difficulty of the regression task, where appropriate distribution assumptions can enhance regression accuracy (as demonstrated by extensive existing studies (Ferrari and Cribari-Neto, 2004; Ouma et al., 2016; Wang et al., 2019a; Mardalena et al., 2020; Shono, 2008)). However, the distribution of short-video watch time remains insufficiently explored. To bridge this knowledge gap, we thoroughly investigate the distribution of short-video watch time using our real-world industrial recommendation data. As shown in Figure 1, we analyze and compare watch time distributions across coarse-to-fine-grained levels:
@@
To mitigate the two challenges mentioned above, and inspired by previous works in other fields (Ferrari and Cribari-Neto, 2004; Ouma et al., 2016; Wang et al., 2019a), we propose to directly regress the absolute values of watch time based on a sound distribution assumption, thereby improving estimation accuracy. To reconcile distributional discrepancies across different granularities, we assume that short-video watch time is governed by a latent mixture model, formally defined as the Exponential-Gaussian Mixture distribution (EGM distribution), where the exponential component addresses coarse-grained distribution skewness and the Gaussian components adaptively capture fine-grained distribution diversity.
@@ 4.2. Main Results
Watch Time Video Views Engage Actions MAE XAUC KL
@@ 4.2.1. Online A/B tests(RQ1)
To further demonstrate the effectiveness of EGMN, we conducted online A/B tests for 7 days on the industrial short video recommender system of Xiaohongshu App. We simultaneously attribute 10% real traffic to EGMN and our existing baseline CREAD. Our platform serves hundreds of millions of DAUs. Thus each traffic group covers enough real users and ensures the reliability of the experimental results. EGMN and CREAD are applied to the ranking stage whose predicting backbone is MMOE (Ma et al., 2018; Tang et al., 2020). We show the results of online A/B testing in the left part of Table 2. It is apparent that the proposed EGMN model achieves a significant improvement (+0.681%) in user watch time compared to the CREAD, while exhibiting no negative impact on other metrics. In addition, it is observed that EGMN also demonstrates a significant improvement (+0.189%) in Video Views, which can be attributed to its better alignment with user interests and ultimate enhancement of user engagement.